Determine the unknown rate. Distance is 500 feet , the time is 9/2 seconds determine the rate please workout problem so that I can understand how to work the problem
Answer by Alan3354(69443)

Determine the unknown rate. Distance is 500 feet , the time is 9/2 seconds determine the rate please workout problem so that I can understand how to work the problem
-----------
D = R*T
R = D/T = 500 ft/4.5 sec
R =~ 111.111... ft/sec